What to do? I don't use calculators and prefer to have written solutions, so HOW?
Solve:
3^x = 10
and
(1.05)^x = 2

Answer by Alan3354(69443)

What to do? I don't use calculators and prefer to have written solutions, so HOW?
Solve:
3^x = 10
x*log(3) = log(10) = 1
x = 1/log(3)
---
You can use a calculator, a slide rule, or a log table.
If you don't want to use a log table, then you'll have to generate your own logs from a series. Are you making some sort of protest?
-----------------
and
(1.05)^x = 2
x*log(1.05) = log(2)
x = log(2)/log(1.05)
-----------------